OpenStudy (anonymous):

how to find the derivative of Ln (x+1/x-1)

OpenStudy (anonymous):

\[f(x)=\ln \frac{x+1}{x-1}\]ok we have a formula like this :\[(\ln u)'=\frac{u'}{u}\]and u have\[u=\frac{x+1}{x-1}\]In order to evaluate \(u'\) use quotient rule